# Compactifications of phylogenetic systems and electrical networks

### v2

@@ 5.2. Corollaries
Counting the compactified split systems for a given circular order on $[n]$ yields
@@
For comparison, the numbers of unweighted equivalence classes of cactus networks for a given circular order are given by $(2n-1)!!$!, which counts the perfect matchings on $[2n]$ For instance, for a given circular order of $[4]$, there are 15 compactified split systems not corresponding to a cactus network — they correspond instead to non-planar cactus networks. At the same time, there are 4 types of unweighted planar cactus networks on $[4]$ that all map to the same split system: the top dimensional cell plus 3 more planar cactus networks whose Kron reduction is the complete graph, as seen in Figure 3. Thus the counts of cells compare: 117 - 15 + 3 = 105.
Moreover, we can also count the compactified electrical networks, using the same arguments as in the proof of Theorem 17, or via a comment of P. Hanna in entry [A111088] of [oeis]:
@@
Counting the compactified electrical networks yields
@@ 5.3. Plabic Tilings and Ranges
Counting phylogenetic networks is an active area of research, with the application of bounding the search time for reconstruction algorithms [4, 23]. An especially useful sort of phylogenetic split system is the induced split system $\mathbf{\sigma}(N)$, as described combinatorially by Gambette, Huber, and Scholz [22]. These are somewhat difficult to count, especially when we allow for multiple connected components, for two reasons. First, the induced split system for networks with multiple connected components is a compactified split system, as first defined in this paper. Second, the fact that the induced map takes disconnected networks to compact split systems, and compact networks to disconnected split systems, means that it is hard to count the split systems in its range using a recursive species procedure, at first glance.
@@
This result points out that the construction of $\tau(N)$ is always possible when $N$ is planar. Thus there are certain obstructions to planarity easily visible in the Kron reduction. Figure 14 shows two networks $N^{\prime}$ and $N^{\prime\prime}$ obtained by adding a wire to network $N$ from Figure 1. The graph of $K(N^{\prime})$ cannot be transformed to a plabic tiling, and therefore $N^{\prime}$ cannot be planar, from Corollary 21. On the other hand, $N^{\prime\prime}$ and $K(N^{\prime\prime})$ show that the converse is not always true: there might be values of the conductances in $N^{\prime\prime}$ making the network planar.
@@
From Theorem 1, and noting that taking the dual is a bijective operation on the (clockwise ordered) compactified electrical networks $\mathbf{\Omega}_{n}$, we conclude that the image of $\mathbf{\Omega}_{n}$ under $\mathbf{\xi}$ is the same as its image under $\mathbf{\sigma}$. We call that image the faithful split systems as in [15]. Thus for compact clockwise networks, the faithful (unweighted) split systems (that is, the unweighted induced systems) are counted by $|\mathbf{\sigma}(\mathbf{\Omega}_{n})|=|\mathbf{\xi}({\mathbf{\Omega}}_{n})|$. These cells form a subcomplex of $\mathbf{\Psi}_{n}.$ The $f$-vector of that subcomplex for $n=4$ is $\langle 14,28,28,20,9,2,1\rangle$.
The total number of (noncompact) unweighted split systems in the image $\mathbf{\xi}(\mathbf{\Omega}_{n})$ in $\mathbf{\Psi}_{n}$, which form a subcomplex of $\mathbf{\Psi}_{n},$ can be counted using a formula found in [26]. There, the Ptolemy diagrams are described as polygons with a subset of diagonals, obeying the rule that for every pair of crossing diagonals in a diagram, all the other diagonals using their four endpoints (edges of a quadrilateral) must be included in that diagram. [^12] The Ptolemy diagrams correspond to torsion pairs in the cluster category of type $A_{n}$ [26]. This rule precisely describes the Kron reductions of circular planar electrical networks (in clockwise order), with the extra requirement that the boundary edges (between consecutive nodes on the boundary) all be included. The Ptolemy diagrams $\mathcal{P}_{n}$ of the $n$-gon for $n=k+3$ are counted in [26] by
[…]
resulting in the OEIS sequence [A181517]. Note that the plabic tilings of a single polygon with $n$ sides are in direct bijection with the Ptolemy diagrams of size $n$, for $n\geq 3$. We also include the unique diagrams of size $n=2$, 1, and 0: a single edge (a trivial split), a single vertex, and the empty diagram.
We wish to count the cells in the image $\mathbf{\xi}(\mathbf{\Omega}_{n}).$ The only difference between the polygonal pictures of split systems in the image of $\mathbf{\xi}$ and the Ptolemy diagrams is that the boundary edges (trivial splits) of the former can be included or excluded, when they are not part of a clique of size 4 or more. Denoting the number of these optional trivial splits for a given diagram $s$ by $t(s)$ we have:
$|\mathbf{\xi}(\mathbf{\Omega}_{n})|\ =\ \sum_{s\in\mathcal{P}_{n}}2^{t(s)}.$

### v3

@@ 5.2. Corollaries
Counting the cactus split systems for a given cyclic order on $[n]$ yields
@@
Some numbers are shown in Table 1. For comparison, the numbers of unweighted equivalence classes of cactus networks for a given cyclic order are given by $(2n-1)!!$!, which counts the perfect matchings on $[2n]$ For instance, for a given cyclic order for $n=4$, there are 15 cactus split systems not corresponding to a cactus circular planar electrical network — they correspond instead to non-planar cactus networks. At the same time, there are 4 types of unweighted planar cactus networks on $[4]$ that all map to the same split system: the top dimensional cell plus 3 more planar cactus networks whose Kron reduction is the complete graph, as seen in Figure 4. Thus the counts of cells compare: 117 - 15 + 3 = 105.
Moreover, we can also count the cactus electrical networks, using the same arguments as in the proof of Theorem 17, or via a comment of P. Hanna in entry [A111088] of [37]:
@@
Counting the cactus electrical networks yields
@@ 5.3. Plabic Tilings and Ranges
Counting phylogenetic networks is an active area of research, with the application of bounding the search time for reconstruction algorithms [4, 22]. An especially useful sort of phylogenetic split system is the induced split system $\mathbf{\sigma}(N)$, as described combinatorially by Gambette, Huber, and Scholz [21]. These are somewhat difficult to count, especially when we allow for multiple connected components, for two reasons. First, the induced split system for networks with multiple connected components is a cactus split system, as first defined in this paper. Second, the fact that the induced map takes disconnected networks to compact split systems, and compact networks to disconnected split systems, means that it is hard to count the split systems in its range using a recursive species procedure, at first glance.
@@
From Theorem 1, and noting that taking the dual is a bijective operation on the (clockwise ordered) cactus electrical networks $\mathbf{\Omega}_{n}$, we conclude that the image of $\mathbf{\Omega}_{n}$ under $\mathbf{\xi}$ is the same as its image under $\mathbf{\sigma}$. We call that image the faithful split systems as in [14]. Thus for compact clockwise networks, the faithful (unweighted) split systems (that is, the unweighted induced systems) are counted by $|\mathbf{\sigma}(\mathbf{\Omega}_{n})|=|\mathbf{\xi}({\mathbf{\Omega}}_{n})|$. These cells form a subcomplex of $\mathbf{\Psi}_{n}.$ The $f$-vector of that subcomplex for $n=4$ is $(14,28,28,20,9,2,1)$.
The total number of (noncompact) unweighted split systems in the image $\mathbf{\xi}(\mathbf{\Omega}_{n})$ in $\mathbf{\Psi}_{n}$, which form a subcomplex of $\mathbf{\Psi}_{n},$ can be counted using a formula found in [26]. There, the Ptolemy diagrams are described as polygons with a subset of diagonals, obeying the rule that for every pair of crossing diagonals in a diagram, all the other diagonals using their four endpoints (edges of a quadrilateral) must be included in that diagram. [^11] The Ptolemy diagrams correspond to torsion pairs in the cluster category of type $A_{n}$ [26]. This rule precisely describes the Kron reductions of circular planar electrical networks (in clockwise order), with the extra requirement that the boundary edges (between consecutive nodes on the boundary) all be included. The Ptolemy diagrams $\mathcal{P}_{n}$ of the $n$-gon for $n\geq 3$, and using $k=n-3$ are counted in [26] by
[…]
resulting in the OEIS sequence [A181517]. Note that the plabic tilings $\mathcal{T}_{n}$ of a single polygon with $n$ sides are in direct bijection with the Ptolemy diagrams of size $n$, for $n\geq 3$. We also include in $\mathcal{T}_{n}$ the unique diagrams of size $n=2$, 1, and 0: a single edge (a trivial split), a single vertex, and the empty diagram. The numbers of these single polygon tilings are then $|\mathcal{T}_{n}|=1,1,1,1,4,17,82,422,...$.
We wish to count the cells in the image $\mathbf{\xi}(\mathbf{\Omega}_{n}).$ The only difference between the polygonal pictures of split systems in the image of $\mathbf{\xi}$ and the Ptolemy diagrams is that the boundary edges (trivial splits) of the former can be included or excluded, when they are not part of a clique of size 4 or more. In terms of our plabic tilings of a polygon, the trivial splits are all boundary edges of a shaded region (or the single edge diagram with no bounded region). Denoting the number of these optional trivial splits for a given plabic tiling $s\in\mathcal{T}_{n}$ by $t(s)$ we have:
Corollary 22.
$|\mathbf{\xi}(\mathbf{\Omega}_{n})|\ =\ \sum_{s\in{\mathcal{T}}_{n}}2^{t(s)}.$
Example.
When $n=4$, the four plabic tilings of a square are $\mathcal{T}_{4}=$[Uncaptioned image]. Thus we have $|\mathbf{\xi}(\mathbf{\Omega}_{4})|=2^{0}+2^{4}+2^{4}+2^{4}=49,$ as seen in Figure 4, inside the dashed line. For $n=0,\dots,6$ we have $|\mathbf{\xi}(\mathbf{\Omega}_{n})|=1,1,2,8,49,373,3196$ respectively.
The proof of Theorem 17 can now be easily used to to count the induced images of the cells of cactus electrical networks inside the cactus circular split systems. Since finding the dual network is a bijection on the cells of $\mathbf{\Omega}_{n},$ then using our Theorem 1, we have that the set of cells of cactus induced systems is the same as the set of cells of cactus graphical systems (images of the graphical map). The important feature is that the cactus graphical systems as a species $A$ are organized precisely as $A=B\circ(X\cdot A)$ where $B$ is the species of ordinary graphical systems. Thus the proof of Theorem 17 gives us:
Corollary 23.
The enumeration of the images of cactus electrical networks yields
$\ |\mathbf{\sigma}(\mathbf{\Omega}_{n})|\ =\ |\mathbf{\xi}(\mathbf{\Omega}_{n})|\ =\ \frac{1}{n+1}\left(\ \sum_{j_{0}+\dots+j_{n}=n}\left(\ \prod_{i=0}^{n}|\mathbf{\xi}(\mathbf{\Omega}_{j_{i}})|\right)\right)\,.$
which can be expanded to:
[…]
where $j_{0},\dots,j_{n}$ are ordered non-negative integers, ${\mathcal{T}}_{j_{i}}$ is the set of plabic tilings of a polygon with $j_{i}$ boundary edges, and $t(s)$ is the number of boundary edges of the shaded regions of $s$.
Some numerical results are seen in Table 1. For example, when $n=3,$ we have the total: $(1/4)(4(8)+12(2)+4(1))=15,$ as pictured in Figure 9. When $n=4,$ we have the total: $(1/5)(5(49)+20(8)+10(4)+30(2)+5(1))=102.$ We would like to find a more efficient formula, but will leave that as an open question.
6. Global Compactified Spaces
6.1.
With future work in mind, we need to extend the definition of the space of circular electrical networks and their compactifications to encompass all circular permutations of $[n]$. We call this extending to the global case. Here we just introduce the spaces and enumerate some cells for small dimensions. Maps of this paper extend to the global spaces : but only $\mathbf{\sigma}$ and $\rho$ are well defined on the new global equivalence classes. The map $\mathbf{\xi}$ requires all its inputs and output to have a fixed cyclic ordering of nodes, so it is only well defined within a single chamber of the global space.
Definition 24.
A global compact circular planar electrical network $N$ is a circular (cactus) network with $n$ boundary nodes labeled by $[n]$, not necessarily in clockwise (nor any specific) order.
Definition 25.
Two global compact circular planar electrical network $N$ and $N^{\prime}$ are equivalent if they both have the same set of pairwise effective resistances $W_{ij}$ between nodes labeled $i$ and $j.$
Note that for instance this means that if $N\equiv N^{\prime}$ they must both have the same sets of identified nodes (where resistance is 0) and connected components, since resistance between disconnected nodes is $\infty$.∞ . Indeed, two planar electrical networks with different cyclic orders $c$ and $c^{\prime}$ of boundary nodes are still equivalent if they have the same response matrix with rows and columns ordered by the usual counting order of $[n].$ Notice that planarity must hold with respect to the cyclic orders: so the two equivalent networks are both planar with respect to both orders.
For any fixed cyclic order $c$ of $[n]$ there is a copy of $\mathbf{\Omega}_{n}$ and of $\mathbf{\Omega}_{n}$, called $\mathbf{\Omega}_{n}^{c}$ and $\mathbf{\Omega}_{n}^{c}$, the circular networks with boundary nodes in the order $c$. Thus the following:
Definition 26.
We define the space $\mathbf{\Omega}^{global}_{n}$ as the union of the $(n-1)!/2$ copies of $\mathbf{\Omega}^{c}_{n}$, called chambers, one for each cyclic order $c$; glued together along equivalent networks. We define the space $\mathbf{\Omega}^{global}_{n}$ as the union of the $(n-1)!/2$ copies of $\mathbf{\Omega}^{c}_{n}$, glued together along equivalent networks.
Similarly we define the global space of circular split systems and its compactification. As mentioned above, a (global) circular split system using any cyclic ordering is actually the sort of circular split system that is usually studied in the phylogenetics literature. Indeed, the Kalmanson (matrix) metrics are defined as having an existing cyclic order for which they obey the conditions, and thus correspond to a system of splits for which there exists a cyclic order of $[n]$ that allows that system to be drawn as a polygonal diagram.
Definition 27.
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on $[n]$, which bijectively correspond to circular split systems.
In [11] these systems are studied with the additional requirement that all trivial splits are present, which reflects a phylogenetic situation in which all taxa are temporally removed from common ancestors. That source points out that visually the diagrams of the split system can be twisted along splits that are either not crossed by another split or not present in the system—and still represent the same split system. When we compactify the space of split systems with a given cyclic order we represented the new points as non-crossing partitions of $[n]$, with the splits on each part of the partition. Now however we allow any partition of $[n]$, and any cyclic order for the polygon on each part of the partition.
Definition 28.
Let $\mathbf{\Psi}^{global}_{n}$ denote the space of Kalmanson metrics on parts of partitions of $[n]$, which correspond bijectively to circular split systems on each part of each partition.
Unweighted cactus networks with any cyclic order correspond to the cells of the resulting CW-complex structure on $\mathbf{\Omega}^{global}_{n}$. The 1-skeletons of $\mathbf{\Omega}^{global}_{n}$ and $\mathbf{\Psi}^{global}_{n}$ are made of 0-cells that have no edge (respectively no split) and 1-cells that have a single edge (single split). Thus these are identical as graphs. Figure 16 shows the 1-skeleton of $\mathbf{\Omega}^{global}_{4}$. Figure 17 shows the 1-skeleton of $\mathbf{\Psi}^{global}_{4}$, and is drawn as the mirror image of Figure 16: the map $\mathbf{\sigma}$ takes electrical networks in Figure 16 to the corresponding split system in the mirror image. Of course the number of higher dimensional cells in $\mathbf{\Omega}^{global}_{n}$ is generally fewer than in $\mathbf{\Psi}^{global}_{n}$ since planarity of the circuit is required, as opposed to just planarity of the splits. The $f$-vector for $\mathbf{\Omega}^{global}_{4}$ is (15, 31, 33, 26, 16, 9, 3) with 133 total cells. The $f$-vector for $\mathbf{\Omega}^{global}_{5}$ is (52, 160, 270, 345, 375, 378, 340, 255, 150, 60, 12) with 2397 total cells. The general numbers of cells for these global spaces, counted here by hand, is an open question, and the first step in an exciting new study of the topology of the spaces themselves. In Table 3 we list some statistics of the examples of spaces in this paper, with OEIS entry names for the sequences that have proven formulas for all $n$. Certain sequences begin with the same terms, since the spaces coincide (and have only one chamber) for $n\leq 3$.
The sequence of numbers of 0-cells for $\mathbf{\Omega}^{c}_{n}$ is shown to be the Catalan numbers in [18]. The sequence of numbers of 0-cells for $\mathbf{\Psi}_{n}$ is clearly the Bell numbers, since those give the numbers of partitions of $n$. That in turn implies that the Bell numbers also count the numbers of 0-cells in $\mathbf{\Omega}_{n}^{global}.$
Figure 16. The 1-skeleton of $\mathbf{\Omega}^{global}_{4}$.
Figure 17. The 1-skeleton of $\mathbf{\Psi}^{global}_{4}$.
Space number of 0-cells total cells $f$-vector
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}^{global}_{2}$ 1 2 (1, 1)
$\mathbf{\Omega}_{2}=\mathbf{\Omega}_{2}^{global}=\mathbf{\Psi}_{2}=\mathbf{\Psi}_{2}^{global}$ 2 3 (2, 1)
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 1 8 (1, 3, 3, 1)
$\mathbf{\Omega}_{3}=\mathbf{\Omega}_{3}^{global}=\mathbf{\Psi}_{3}=\mathbf{\Psi}^{global}_{3}$ 5 15 (5, 6, 3, 1)
$\mathbf{\Omega}_{4}$ 1 52 [A111088] (1, 6, 14, 16, 10, 4, 1)
$\mathbf{\xi}(\mathbf{\Omega}_{4})$ 1 49 [Cor. 22] (1, 6, 14, 16, 9, 2, 1)
$\mathbf{\Psi}_{4}$ 1 64 [A006125] (1, 6, 15, 20, 15, 6, 1)
$\mathbf{\Omega}_{4}$ 14 [A000108] 105 [A001147] (14, 28, 28, 20, 10, 4, 1)
$\mathbf{\xi}(\mathbf{\Omega}_{4})=\mathbf{\sigma}(\mathbf{\Omega}_{4})$ 14 [A000108] 102 [A136654] (14, 28, 28, 20, 9, 2, 1)
$\mathbf{\Psi}_{4}$ 14 [A000108] 117 [A136654] (14, 28, 29, 24, 15, 6, 1)
$\mathbf{\Omega}^{global}_{4}$ 1 70 [open] (1, 6, 15, 20, 16, 9, 3)
$\mathbf{\Psi}_{4}^{global}$ 1 112 [open] (1, 7, 21, 34, 31, 15, 3)
$\mathbf{\Omega}^{global}_{4}$ 15 [A000110] 133 [open] (15, 31, 33, 26, 16, 9, 3)
$\mathbf{\sigma}(\mathbf{\Omega}^{global}_{4})$ 15 [A000110] 124 [open] (15, 31, 33, 26, 13, 3, 3)
$\mathbf{\Psi}_{4}^{global}$ 15 [A000110] 169 [open] (15, 31, 36, 38, 31, 15, 3)
Table 3. CW-complexes in this paper: for $n=2,3,4$, some sequences and $f$-vectors.
The (known terms in) the sequence of total numbers of cells of $\mathbf{\Psi}^{global}_{n}$ are 2, 8, 112, 6976, 1332224. These are based on the hand-counted values in [39], each multiplied by $2^{n}$ to allow any subset of the trivial splits to be included in the split system. The general term is an open question.
7. Consequences
7.1. Corollaries
Here we see several implications of the main theorems. Finding the split system $\rho(N)=\mathbf{\sigma}(N)$ in Theorem 8 is our contribution to Question (1) from the Introduction. In [13], we showed that the large-scale, galled-tree, structure of the network $N$ is captured by that of the split system $\rho(N)$. The algorithm Neighbor-net takes any resistance matrix (perhaps of a circular electrical network), without regard for cyclic order, and returns a split network that exhibits a cyclic order and which reproduces precisely the tree-like portions of the circular planar electrical network. Recall that a bridge is a split that is drawn as a diagonal that does not cross any others. The following corollary to Theorem 8 is proven in [13]:
Corollary 29.
The bridges and cut vertices of $N$ all become bridges in $\rho(N).$ Specifically, the bridges of $N$ are found among the bridges of $\rho(N).$
This corollary allows the problem of reconstructing a network from its response matrix to be subdivided into constructing sub-networks that are separated by bridges. The bridges are made visible in the split system. This preservation of bridges allows the modular reconstruction of $N$ from its response matrix, as demonstrated in [13]. We next discuss the question of which cyclic orders can be consistent with a given network.
Definition 30.
For an $n$-node circular planar electrical network $N$, a cyclic order $c$ of $[n]$ is consistent with $N$ when a new network can be created with that cyclic order of terminal nodes, preserving both the planarity and the pairwise resistances of $N.$ The set of consistent cyclic orders is denoted $\mathcal{O}_{c}(N)$.
Just looking at a diagram of a circular electrical network with crossing wires, it can be hard to determine whether it is in fact planar. One nice feature of finding its split system $\rho(N)=\mathbf{\xi}^{\prime}(N*)$ is that if $N$ is actually planar, then its set of cyclic orders will be visible in the picture of the split system. That visibility follows from the fact that the split system is unique, and that two pictures of the split system (with different cyclic orders) can always be transformed into each other via the twists along diagonals described in [11]. We make this more precise with our newly defined plabic tiling: $\tau(N)$ is thus useful for enumerating the consistent cyclic orders $\mathcal{O}_{c}(N)$, that is the number of top-dimensional cells which contain the cell of $N.$ We show that the number of cyclic orders of $[n]$ consistent with $N$ is found by considering the sizes of the regions of $\tau(N),$ answering Question (3) from the Introduction.
Theorem 31.
The number of cyclic orders consistent with a given connected, single-bulbed, circular planar electrical network $N$ is found by one half the product of factors, one factor for each region of $\tau(N).$ Each shaded region in $\tau(N)$ with $k$ sides contributes a factor of $(k-1)!$ and each unshaded region contributes a factor of 2.
Proof.
Since $N$ is planar, $\tau(N)$ will reflect the structure of the unique split system $\rho(N),$ which we can draw as a split network. Shaded regions correspond to cut-nodes of the split network: $k$ portions of the network meet at that cut node and thus can be arranged in any cyclic order, contributing $(k-1)!$ options. The unshaded regions can be flipped over, so each contributes 2 more independent options. Finally choosing various of these options can eventually flip over the entire network which is really the same cyclic ordering, so we need to divide by 2 in the end. ∎
Example.
The plabic tiling $\tau(N)$ in Figure 15 has 3 shaded regions contributing factors of $(3-1)!,(3-1)!$ and $(4-1)!$ respectively, and one unshaded region contributing a factor of 2. By Theorem 31, the number of consistent cyclic orders for the network $N$ from Figure 2 is $|\mathcal{O}_{c}(N)|=2(2!)(2!)(3!)/2=24.$
The cases of a disconnected or cactus network with more than one bulb are more complicated, and we leave those for future work. For now we mention a consequence that answers a question which is perhaps already familiar to circuit designers: Rearranging the terminal nodes of a planar network might cause wires to cross. And if that new cyclic order is not consistent with the split system, it will not allow the new arrangement to be equivalent to a planar one.
Corollary 32.
Consider $N^{\prime}$ formed by rearranging the boundary nodes of $N$ while keeping all edges intact, but in a cyclic order $c^{\prime}$ not consistent with $\rho(N)$. Then $N^{\prime}$ with its new cyclic order $c^{\prime}$ is guaranteed to be nonplanar.
This is exemplified in Figure 18, where $N^{\prime}$ is obtained from $N$ by transposing terminals $5$ and $6$. To see that the resulting $N^{\prime}$ is non-planar, we can check the absence of its cyclic order from the list of 24 consistent circular orders $\mathcal{O}_{c}(N).$ Notice that the graphs of $K(N)$ and $K(N^{\prime})$ are the same shape; but the response matrices using the two orders to order their rows would have differing values, and their minors would reflect the planarity and non-planarity respectively. Corollary 32 simply allows that knowledge without directly calculating the minors.
Figure 18. The graph of $N^{\prime}$ cannot be planar, despite sharing $K(N)\cong K(N^{\prime})$.
Corollary 21 points out that the construction of $\tau(N)$ is always possible when $N$ is planar. That fact allows our answer to Question (2) in the Introduction: there are certain obstructions to planarity easily visible in the Kron reduction.
Example.
@@ 7.1. Corollaries
Figure 19 shows two networks $N^{\prime}$ and $N^{\prime\prime}$ obtained by adding a wire to network $N$ from Figure 2. The graph of $K(N^{\prime})$ cannot be transformed to a plabic tiling, and therefore $N^{\prime}$ cannot be planar, from Corollary 21. On the other hand, $N^{\prime\prime}$ and $K(N^{\prime\prime})$ show that the converse is not always true.
@@
7.2. Traveling Salesman Polytopes
Since the Kalmanson and cographical split systems coincide, we can transfer known results relating split systems and the Symmetric raveling Salesman Polytope, to the new arena of circular planar electrical networks.
Definition 33.
For a cyclic order $c$ of $[n]$ let the incidence vector $\mathbf{x}(c)$ have components for each pair $i<j\in[n]$; equaling 1 if $i,j$ are adjacent in $c$, 0 if not. The ordering of the components is lexicographic. These incidence vectors for all cyclic orders make up the $(n-1)!/2$ vertices of the Symmetric Traveling Salesman Polytope STSP($n$).
Definition 34.
For a given circular planar electrical network $N$ define by $W_{N}(\mathbf{x})=W_{N}\cdot\mathbf{x}$ the linear functional where the dot product is taken with the upper triangle of $W(N)$, not including the diagonal, read by rows.
Theorem 35.
For a circular planar electrical network $N$, the linear functional $W_{N}$ is minimized simultaneously over the STSP($n$) at the vertices $\mathbf{x}(c)$ corresponding to the orders in $\mathcal{O}_{c}(N)$ The value of that minimum is twice the sum of the splits of $\rho(N).$
Proof.
This follows from Theorem 5.6 of [15] and Theorem 4.5 of [14]. The key is that any circuit of the boundary nodes in consistent cyclic order will accumulate two copies of every split. ∎